if xy=1 then y^2 +dy/dx is =??
1
Expert's answer
2012-10-09T11:32:48-0400
xy=1
y=1/x

dy/dx=-1/x^2

-1/x^2=-y^2

y^2-dy/dx=y^2-y^2=0